OpenStudy (anonymous):

Decide whether the pair of lines is parallel, perpendicular, or neither. 3x - 8y = 8 and 32x + 12y = 2

OpenStudy (aroub):

-8y=-3x+8 y=3/8-1 12y=-32x+2 y=-8/3 + 1/6 so they are perpendicular.